OpenStudy (anonymous):

solve : x − 8 = √(-18x + 67)

OpenStudy (anonymous):

start with \((x-8)^2=-18x+67\) solve the quadratic equation and don't forget to check your solution in the original equation

OpenStudy (anonymous):

1 or -3

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

x^2+64-16x=-18x+67 x^2+2x-3=0 (x-1)(x+3)=0

OpenStudy (anonymous):

so there r 2 solutions

OpenStudy (anonymous):

hmmm really?

OpenStudy (anonymous):

if you square both sides of an equation, you can introduce extraneous solutions we know \(-5\) is not a solution to \(x+1=4\) but if we square both sides, it is a solution to \((x+1)^2=16\) by squaring, we have introduces an additional solution, namely \(-5\) to the original equation \(x+1=4\)

OpenStudy (anonymous):

that is why you have to check both solutions after you square, to see if they work. \(-3\) is not a solution to \[x-8=\sqrt{-18x+67}\]